OpenStudy (anonymous):

please help :) A particle is moving along the curve y=5sqrt(4x+4). As the particle passes through the point (3,20) its x-coordinate increases at a rate of 3 units per second. Find the rate of change of the distance from the particle to the origin at this instant.

OpenStudy (anonymous):

I have... dy/dt=(3)(0.3)[(4x+4)^(-1/2)](5dx/dt)

OpenStudy (anonymous):

dy/dt = 4(0.5)[(4x+9)^(-1/2)](4dx/dt) so when x=4 dy/dt = (8/5)(dx/dt)=8 units/sec D = (x^2+y^2)^0.5 dD/dt = 0.5(x^2+y^2)^(-0.5) (2dx/dt+2ydy/dt) = (5 + 20*8) / sqrt(16+400) = 165/20.396=8.089 units/sec

OpenStudy (anonymous):

that's not right.. that's what I had the first time and it said it was wrong.. I don't know what I'm doing wrong.

OpenStudy (anonymous):

@Chlorophyll can you help me? I have done this problem over and over and can't seem to get the right answer?

OpenStudy (anonymous):

guess you need the distance formula right?

OpenStudy (anonymous):

I don't know what I need.. I obviously can't seem to get it right... It's bothering me :/

OpenStudy (anonymous):

or at least the square of the distance square of the distance from \((0,0)\) so \((x,5\sqrt{4x+4})\) is \[d^2=x^2+25(4x+4)\]

OpenStudy (anonymous):

sqrt(409) right?

OpenStudy (anonymous):

we can simplify this a little \[d^2=x^2+100x+100\] and you are looking for \(d'\) taking derivatives wrt \(t\) using the chain rule we get \[2dd'=2xx'+100x'\]

OpenStudy (anonymous):

maybe easier to start with \[dd'=xx'+100x'\] and you are told that \(x'=3\) and also that \(x=3\) so you can solve for \(d'\)

OpenStudy (anonymous):

oh damn i mean \[dd'=xx'+50x\] sorry

OpenStudy (anonymous):

dd'=159?

OpenStudy (anonymous):

now \(x=3,x'=3\) so we get \[dd'=9+150=159\] right

OpenStudy (anonymous):

you want \(d'\) so all that is left is to find \(d\) which you also know since it is the distance between \((0,0)\) and \((3,20)\)

OpenStudy (anonymous):

i get \(d=\sqrt{409}\) making \(d'=\frac{159}{\sqrt{409}}\) but my arithmetic sucks, so i wouldn't bet money on it

OpenStudy (anonymous):

that is the idea however

OpenStudy (anonymous):

I got sqrt(409) as well.

OpenStudy (anonymous):

ok then i am happy with that answer

OpenStudy (anonymous):

you?

OpenStudy (anonymous):

my gosh, thank so much...

OpenStudy (anonymous):

yeah, it's correct!

OpenStudy (anonymous):

yw, just hope it is right notice, however, that when working with distance problems it is much easier to work with the square of the distance makes all the computations a lot nicer